Find the equation to the straight line which passes through the point (4,-5) and which is parallel to straight line (3x+4y+5=0)

Grade:10

1 Answers

Aditya Gupta
4 years ago
slope of given line 3x+4y+5=0 is – ¾
so any line parallel to it will also have the same slope.
by point slope form, we write eqn of reqd line as:
y+5= – ¾ *(x – 4)
or 3x+4y+8= 0
samaira please approve yar. :=))